# Faithful Logic Embeddings in HOL — Deep and Shallow

### v2

@@ 6 Minimal shallow embedding of PML in HOL
To implement this idea, ${W}$ is now implicitly identified with type w of HOL at the metalevel, and the constant symbols R and V are introduced in metalogic HOL to represent the semantic notions ${R}$ and ${V}$ of PML (see line 5). Most importantly, the modified HOL type $\sigma:=\texttt{w}{\Rightarrow}\texttt{bool}$ is now taken as the type of the embedded formulas $\varphi_{\sigma}$, indicating that PML formulas in this minimal embedding now depend only on possible worlds. [^5] Note that this is also the core idea of the well-known standard translation of PML into first-order logic [18], which, in fact, is implemented here by exploiting $\lambda$-conversion and compositionality directly in the metalogical HOL without the need for explicit recursive definitions. Hence, instead of maintaining dependencies on ${W}$, ${R}$, ${V}$, and $w$, the definitions of the logical connectives $a^{m}$, $\neg^{m}$, $\supset^{m}$, and $\Box^{m}$ now only explicitly maintain a dependency on worlds $w$ (lines 8-11). Their definitions are thus, where V and R are the above-mentioned constant symbols in metalogic HOL: $a^{s}:=\lambda w.\,\texttt{V}\,a\,w$, $\neg^{m}\varphi:=\lambda w.\,\neg(\varphi\,w)$ and $\varphi\supset^{m}\psi:=\lambda w.\,\varphi\,w\longrightarrow\psi\,w$, and $\Box^{m}\varphi:=\lambda w.\,\forall v.\,\texttt{R}\,w\,v\longrightarrow\varphi\,v$. The definition of the other logical connectives $\vee^{m}$, $\wedge^{m}$, $\Diamond^{m}$, $\top^{m}$, and $\bot^{m}$ (in lines 13-17) is analogous to before. Truth in a world $w$, for a given model $\langle{W},{R},{V}\rangle$ maintained at the meta-lavel HOL, called RelativeTruthM and noted as $w\models^{m}\varphi$, is then defined (see line 19) as the application of the embedded formula $\varphi$ to $w$. Consequently, validity $\models^{m}\varphi$ of an embedded formula $\varphi$ is defined as relative truth w.r.t. to all worlds $w$ (line 21). The bag DefM defined (in lines 21-23) is analogous to DefD and DefS from before. Again, not a single axiom is introduced. So PML is once again characterized as a proper fragment of HOL, but now more lightweight.
@@ 7 Faithfulness automated
$\forall\,WR\,V\,w.\ \langle W,R,V\rangle,w\models^{d}\varphi$ $\longleftrightarrow\langle W,R,V\rangle,w\models^{s}\llparenthesis\varphi\rrparenthesis$ (Faithful1a)
@@
$\forall\,w.\ \langle(\lambda x_{w}.\top),\texttt{R},\texttt{V}\rangle,w\models^{d}\llparenthesis\varphi\rrparenthesis$ $\longleftrightarrow w\models^{m}\llbracket\varphi\rrbracket$ (Faithful3)
@@ 8 Experiments
The possibilities that the linked embeddings offer, e.g., for logic education, are illustrated with some experiments. The experiments performed also test the soundness of our embeddings, and assess proof automation and counterexample finding for them with sledgehammer and nitpick. The embedding that has performed best in this regard is the minimal shallow embedding. This is why file PMLinHOL_shallow_minimal_tests.thy, see Fig. 4, is presented here. [^6] Corresponding files for the deep and the maximal shallow embeddings are shown in the appendix [5]. The Isabelle/HOL sources are available at http://logikey.org/tree/master/CoursesAndTutorials/2025-CADE-DeepShallow.
@@
The first experiments (see lines 4-7) confirm that the schematic axioms H1-H3 and the modus ponens rule MP of the propositional Hilbert calculus from §2 are implied by the minimal shallow embedding (local’ versions of H1-H3 and MP rule are also implied; this is not shown here). The auto tactic, and with it, the simplifier simp are used, which has access to the definitions in bundle DefM. The $\varphi$, $\psi$ and $\gamma$ are schematic variables of type $\sigma$ in the metalogic HOL. The proofs were originally found by calling sledgehammer.
@@
The next experiment (in lines 9-14) demonstrates automated and interactive reasoning using the schematic axioms and rules of the Hilbert calculus introduced earlier. It is shown that the formula schema $\varphi\supset^{m}\varphi$ is derivable in the given Hilbert calculus. By preventing access to the definitions in the bag DefM, e.g. with the command sledgehammer(H1 H2 H3 MP), the automated theorem provers orchestrated by sledgehammer can be forced to search for proofs in the given Hilbert calculus only, and they succeed here. They find out that only H1 and H2 and MP are needed to obtain the result. In the given situation reconstruction of a proof from H1 and H2 and MP is then possible with Isabelle’s trusted metis tool. For educational purposes it is relevant that proofs from the literature can be easily mapped one to one to proofs in our setting. This is illustrated here with the provision of an interactive proof (in lines 10-14) which is following a proof from a Wikipedia entry [62] on Hilbert systems. Further derived axiom schemata are proved automatically at the level of the Hilbert calculus (in lines 15-16). Of course, such attempts can also fail (as was the case in line 17), and then either interactive proof is needed or, as demonstrated here with the call of auto, using the definitions of the shallow embedding in metalogic HOL may in fact help to still find a proof automatically. Depending on prior knowledge and teaching objectives, students can be encouraged to provide interactive proofs in exercises rather than using automation tools. It is also important to note that Isabelle/HOL’s proof tactics support, and in particular the Eisbach package, can be used to further enhance the automation of proofs at the level of embedded logic. [^7] The Eisbach package was e.g. used in [40, 41] to define and automate proof methods for a non-trivial shallow embedding of abstract object theory AOT [63] in HOL.
@@
Even more interesting than automatically proving the above correspondences is disproving falsely assumed correspondences. For example, when nitpick is called on the false conjecture, that in PML models with reflexive accessibility relations $R$ the axiom schema $\Box^{m}\varphi\supset^{m}\Box^{m}\Box^{m}\varphi$ is implied (see line 28), the counterexample shown in Fig. 5 is reported back.
@@
Figure 6 presents further experiments. After proving the schematic modal principle $K\Diamond$ (in line 5) an object level proof example [55, Ex. 6.10] is studied (in lines 8-16): $\Box^{m}p^{m}\supset^{m}(\Diamond^{m}q^{m}\supset^{m}\Diamond^{m}(p^{m}\wedge^{m}q^{m}))$, where $p$ and $q$ are propositional symbols from the PML signature $\mathcal{S}$. This statement is again effectively automated via definition unfolding at the metalevel HOL (line 8), while the presented interactive proof (in lines 9-16)uses the propositional Hilbert calculus from before in combination with the modal rule of necessitation, the distribution axiom of base logic $\mathbf{K}$ and the newly introduced $K\Diamond$-principle.
@@
Larger object-level proof examples could now be presented and studied using the minimal shallow embedding. Related studies have shown that the this approach is generally competitive, in particular, when it comes to reasoning in quantified modal logics; see e.g. the recent experiments reported in [59]. Recipes for embedding modal quantifiers and extensive application studies for quantified modal logics are available in prior works; see e.g. [12, 16].
@@
The performance of sledgehammer and nitpick for the different embeddings for the example proof problems as discussed in this paper so far is briefly discussed: [^10] A MacBook Pro 16,1 (6-Core Intel Core i7 processor, 2,6 GHz, 6 Cores, 256 KB L2 Cache per Core, 12 MB L3 Cache, 16 GB Memory) and Isabelle 2025 in standard configuration was used, which sets a 30s prover timeout for sledgehammer. Minimal shallow embedding: Proofs for 36 out of 38 valid statements (theorems, lemmas and substatements in interactive proofs) were found by calls to sledgehammer. Nitpick reported counterexamples for 5 out of 5 invalid statements. Reconstruction of sledgehammer proofs with trusted tactics failed in 2 cases. Maximal shallow embedding: Proofs for 34 out of 38 valid statements were foundsledgehammer. Countermodels for 4 out of 5 invalid statements were found by nitpick. Reconstruction of sledgehammer proofs with trusted tactics failed in 4 cases. Deep embedding: Proofs for 32 out of 38 valid statements were found sledgehammer; in particular, there were no proofs delivered for the Löb examples. Countermodels for 5 out of 5 invalid statements were found by nitpick. Reconstruction of proofs failed in 2 cases.
@@
Further experiments were carried out in order to assess the performance of the different embeddings more rigorously. These tests were about the classification of examples of PML formulas with respect to the logics in the modal logic cube [33] defined by the schematic axioms $T:=\Box\varphi\supset\varphi$, $B:=\varphi\supset\Box\Diamond\varphi$ and $4:=\Box\varphi\supset\Box\Box\varphi$, respectively by their semantic counterparts $Ts:=\texttt{reflexiv}\,R$, $Bs:=\texttt{symmetric}\,R$ and $4s:=\texttt{transitiv}\,R$ (with $R$ being the accessibility relation). The example formulas studied were:
@@
Figure 8: Performance results of automated PML formula classification task conducted with nitpick and sledgehammer (explanation: the first entry $73(\textbf{38}/35)$ in column $D$, e.g., states that 72 proofs were reported in total by sledgehammer for the problem set $M\models^{m}F$, with 38 proofs found with prior application of the simplifier simp and 35 without; analogous for the other entries).
@@
The results obtained are briefly discussed. Perfect experimentation results were obtained for the problem set $N\models^{m}F$, that is, the minimal shallow embedding when used in combination with the semantical frame conditions $Ts$, $Bs$ and $4s$. In this setting nitpick either provided a counterexample for each formula classification task or reported that there was no (finite) counterexample. The latter judgments were then supplemented by proofs from sledgehammer that confirmed the validity of these formulas. The second best performance was achieved with the minimal shallow embedding when used in combination with the axiom schemata $T$, $B$ and $4$. The maximal shallow embedding returned some good results regarding counterexample finding with nitpick and theorem proving with sledgehammer for the problem set $N\models^{s}F$ when prior simplification was applied. However, in this setting, nitpick was unable to identify any formulas with no (finite) countermodels. For the deep embedding the performance results obtained with nitpick regarding counterexample finding significantly dropped in comparison to the minimal shallow embedding. The numbers reported for the deep embedding regarding the identification of formulas with no (finite) countermodels are still good though. [^11] However, some seemingly self-contradictory answers by nitpick in this setting with and without prior simplification for some problems require further consideration. Interestingly, the maximal number of 38 proofs was still obtained for by sledgehammer in the deep embedding when prior simplification was applied. Future work should extend the experiments presented to more complicated test formulas $F_{i}$ and also include the further axiom schemata (and their corresponding semantic conditions) from the modal logic cube [33]. Additionally, it seems useful to replay, now using the embeddings provided in this paper, the metalogical, modal logic cube verification experiments reported in prior work [3, 8].
@@ 9 Related work and implications for LogiKEy
The notions of deep and shallow embeddings appear to originate from the work of Boulton et al. [24], and the connections between the two have been discussed in numerous related papers, see e.g. [60, 39, 52, 61] and the references therein. However, to the best of the author’s knowledge, no previous work has achieved an integration of deep and (different degrees of) shallow embeddings in the form presented, enabling in particular mutual faithfulness proofs between them.
@@
The logic-pluralistic knowledge and reasoning methodology LogiKEy[10], see Fig. 9, leverages higher-order logic (HOL) as a unifying metalogic to enable the modeling of diverse object logics, so far, primarily via using shallow semantic embeddings. LogiKEy has been used in a range of application domains and contexts. Very different kinds of object logics have been been embedded LogiKEy over the past decade, including, for example, free logic [14], modal higher-order logic [16], public announcement logic [13], conditional logic [4] and dyadic deontic logic with a preference-based semantics [50].
@@
This paper also illustrates the conceptual modifications made in previous work when moving from lightweight shallow embeddings of non-classical logics, such as those used to encode higher-order modal logic [12] or conditional logic [4], which consider only dependencies on possible worlds, to more heavyweight shallow embeddings, as, for example, required to embed public announcement logic [13]; in the latter work additional dependencies on evaluation domains had to be encoded. Note on the other hand, that previous work has often avoided the explicit modeling of atoms, signatures, and evaluations, which has been essential for the work presented in this paper. Moreover, the presented technique should eventually help to overcome challenges in application contexts that require the combined use of deep and shallow embeddings. The Fatio protocol for multi-agent argumentation [51] is such an example. Most importantly, the presented technique for linking deep and shallow embeddings will be very fruitful for furthering the use of LogiKEy in university-level logic courses, including the author’s courses on universal logical reasoning that have been taught for many years at FU Berlin and, more recently, at the University of Bamberg.

### v3

@@ 6 Minimal shallow embedding of PML in HOL
To implement this idea, ${W}$ is now implicitly identified with type w of HOL at the metalevel, and the constant symbols R and V are introduced in metalogic HOL to represent the semantic notions ${R}$ and ${V}$ of PML (see line 4). Most importantly, the modified HOL type $\sigma:=\texttt{w}{\Rightarrow}\texttt{bool}$ is now taken as the type of the embedded formulas $\varphi_{\sigma}$, indicating that PML formulas in this minimal embedding now depend only on possible worlds. [^5] Note that this is also the core idea of the well-known standard translation of PML into first-order logic [17], which, in fact, is implemented here by exploiting $\lambda$-conversion and compositionality directly in the metalogical HOL without the need for explicit recursive definitions. Hence, instead of maintaining dependencies on ${W}$, ${R}$, ${V}$, and $w$, the definitions of the logical connectives $a^{m}$, $\neg^{m}$, $\supset^{m}$, and $\Box^{m}$ now only explicitly maintain a dependency on worlds $w$ (lines 7-10). Their definitions are thus, where V and R are the above-mentioned constant symbols in metalogic HOL: $a^{s}:=\lambda w.\,\texttt{V}\,a\,w$, $\neg^{m}\varphi:=\lambda w.\,\neg(\varphi\,w)$ and $\varphi\supset^{m}\psi:=\lambda w.\,\varphi\,w\longrightarrow\psi\,w$, and $\Box^{m}\varphi:=\lambda w.\,\forall v.\,\texttt{R}\,w\,v\longrightarrow\varphi\,v$. The definition of the other logical connectives $\vee^{m}$, $\wedge^{m}$, $\Diamond^{m}$, $\top^{m}$, and $\bot^{m}$ (in lines 12-16) is analogous to before. Truth in a world $w$, for a given model $\langle{W},{R},{V}\rangle$ maintained at the meta-lavel HOL, called RelativeTruthM and noted as $w\models^{m}\varphi$, is then defined (see line 18) as the application of the embedded formula $\varphi$ to $w$. Consequently, validity $\models^{m}\varphi$ of an embedded formula $\varphi$ is defined as relative truth w.r.t. to all worlds $w$ (line 20). The bag DefM defined (in lines 22-24) is analogous to DefD and DefS from before. Again, not a single axiom is introduced. So PML is once again characterized as a proper fragment of HOL, but now more lightweight.
@@ 7 Faithfulness automated
$\forall\,WR\,V.\forall w{:}W.\ \langle W,R,V\rangle,w\models^{d}\varphi$ $\longleftrightarrow\langle W,R,V\rangle,w\models^{s}\llparenthesis\varphi\rrparenthesis$ (Faithful1a)
@@
$\forall\,w.\ \langle(\lambda x_{w}.\top),\texttt{R},\texttt{V}\rangle,w\models^{s}\llparenthesis\varphi\rrparenthesis$ $\longleftrightarrow w\models^{m}\llbracket\varphi\rrbracket$ (Faithful3)
@@ 8 Experiments
The possibilities that the linked embeddings offer, e.g., for logic education, are illustrated with some experiments. The experiments performed also test the soundness of our embeddings, and assess proof automation and counterexample finding for them with sledgehammer and nitpick. The embedding that has performed best in this regard is the minimal shallow embedding. This is why file PMLinHOL_shallow_minimal_tests.thy, see Fig. 4, is presented here. [^6] Corresponding files for the deep and the maximal shallow embeddings are shown in the appendix. The Isabelle/HOL sources are available at http://logikey.org/tree/master/CoursesAndTutorials/2025-CADE-DeepShallow.
@@
The first experiments (see lines 4-7) confirm that the schematic axioms H1-H3 and the modus ponens rule MP of the propositional Hilbert calculus from §2 are implied by the minimal shallow embedding (local versions of H1-H3 and MP rule are also implied; this is not shown here). The auto tactic, and with it, the simplifier simp are used, which has access to the definitions in bundle DefM. The $\varphi$, $\psi$ and $\gamma$ are schematic variables of type $\sigma$ in the metalogic HOL. The proofs were originally found by calling sledgehammer.
@@
The next experiment (in lines 9-14) demonstrates automated and interactive reasoning using the schematic axioms and rules of the Hilbert calculus introduced earlier. It is shown that the formula schema $\varphi\supset^{m}\varphi$ is derivable in the given Hilbert calculus. By preventing access to the definitions in the bag DefM, e.g. with the command sledgehammer(H1 H2 H3 MP), the automated theorem provers orchestrated by sledgehammer can be forced to search for proofs in the given Hilbert calculus only, and they succeed here. They find out that only H1 and H2 and MP are needed to obtain the result. In the given situation, the reconstruction of a proof from H1 and H2 and MP is then possible with Isabelle’s trusted metis tool. For educational purposes it is relevant that proofs from the literature can be easily mapped one to one to proofs in our setting. This is illustrated here with an interactive derivation (in lines 10-14) that is following a proof from a Wikipedia entry [61] on Hilbert systems. Further derived axiom schemata are proved automatically at the level of the Hilbert calculus (in lines 15-16). If such attempts fail (as was initially the case in line 17), then either interactive proof is needed or, as is demonstrated here with the call of auto, the unfolding of the definitions of the shallow embedding in metalogic HOL (within auto) may in fact help to still find a proof automatically. Depending on prior knowledge and teaching objectives, students can be asked to solely provide interactive proofs in such exercises rather than using automation tools. It is also important to note that Isabelle/HOL’s proof tactics support, and in particular the Eisbach package, can be used to further enhance the automation of proofs at the level of embedded logic. [^7] The Eisbach package was e.g. used in [39, 40] to define and automate proof methods for a non-trivial shallow embedding of abstract object theory AOT [62] in HOL.
@@
Even more interesting than automatically proving the Sahlqvist correspondences is disproving falsely assumed correspondences. For example, when nitpick is called on the false conjecture, that in PML models with reflexive accessibility relations $R$ the axiom schema $\Box^{m}\varphi\supset^{m}\Box^{m}\Box^{m}\varphi$ is implied (see line 28), the counterexample shown in Fig. 5 is reported back.
@@
Figure 6 presents further experiments. After proving the schematic modal principle $K\Diamond$ (in line 4) an object level proof example [54, Ex. 6.10] is studied (in lines 6-14): $\Box^{m}p^{m}\supset^{m}(\Diamond^{m}q^{m}\supset^{m}\Diamond^{m}(p^{m}\wedge^{m}q^{m}))$, where $p$ and $q$ are propositional symbols from the PML signature $\mathcal{S}$. This statement is again effectively automated via definition unfolding at the metalevel HOL (line 6), while the presented interactive proof (in lines 8-14) uses the propositional Hilbert calculus from before in combination with the modal rule of necessitation, the distribution axiom of base logic $\mathbf{K}$ and the newly introduced $K\Diamond$-principle.
@@
Larger object-level proof examples could now be presented, studied and automated using the minimal shallow embedding. Prior studies have shown that this approach is generally competitive, in particular, when it comes to automating quantified modal logics; see e.g. the recent experiments reported in [58]. Recipes for embedding modal quantifiers and extensive application studies for quantified modal logics have been presented in prior works; see e.g. [11, 15].
@@
The performance of sledgehammer and nitpick for the different embeddings for the example proof problems as discussed so far is briefly discussed: [^10] A MacBook Pro 16,1 (6-Core Intel Core i7 processor, 2,6 GHz, 6 Cores, 256 KB L2 Cache per Core, 12 MB L3 Cache, 16 GB Memory) and Isabelle 2025 in standard configuration was used, which sets a 30s prover timeout for sledgehammer. Minimal shallow embedding: Proofs for 36 out of 38 valid statements (theorems, lemmas and substatements in interactive proofs) were found by calls to sledgehammer. Nitpick reported counterexamples for 5 out of 5 invalid statements. Reconstruction of sledgehammer proofs with trusted tactics failed in 2 cases. Maximal shallow embedding: Proofs for 34 out of 38 valid statements were found by sledgehammer. Countermodels for 4 out of 5 invalid statements were found by nitpick. Reconstruction of sledgehammer proofs with trusted tactics failed in 4 cases. Deep embedding: Proofs for 32 out of 38 valid statements were found by sledgehammer; in particular, there were no proofs delivered for the Löb examples. Countermodels for 5 out of 5 invalid statements were found by nitpick. Reconstruction of proofs failed in 2 cases.
@@
Further experiments were carried out in order to assess the performance of the different embeddings more rigorously. These tests focused on the classification of PML formula examples with respect to the logics in the modal logic cube [32] defined by the schematic axioms $T:=\Box\varphi\supset\varphi$, $B:=\varphi\supset\Box\Diamond\varphi$ and $4:=\Box\varphi\supset\Box\Box\varphi$, respectively by their semantic counterparts $Ts:=\texttt{reflexiv}\,R$, $Bs:=\texttt{symmetric}\,R$ and $4s:=\texttt{transitiv}\,R$ (with $R$ being the accessibility relation). The example formulas studied were:
@@
Figure 8: Performance results of the automated PML formula classification task conducted with nitpick and sledgehammer (explanation: the first entry $73(\textbf{38}/35)$ in column $D$, e.g., states that 73 proofs were reported in total by sledgehammer for the problem set $M\models^{m}F$, with 38 proofs found with prior application of the simplifier simp and 35 without; analogous for the other entries).
@@
The results obtained are briefly discussed. Perfect experimentation results were obtained for the problem set $N\models^{m}F$, that is, the minimal shallow embedding when used in combination with the semantical frame conditions $Ts$, $Bs$ and $4s$. In this setting nitpick either provided a counterexample for each formula classification task or reported that there was no (finite) counterexample. The latter judgments were then supplemented by proofs from sledgehammer that confirmed the validity of these formulas. The second best performance was achieved with the minimal shallow embedding when used in combination with the axiom schemata $T$, $B$ and $4$. The maximal shallow embedding returned some good results for the problem set $N\models^{s}F$ regarding counterexample finding with nitpick and regarding theorem proving with sledgehammer when prior simplification was applied. However, in this setting, nitpick was unable to identify any formulas with no (finite) countermodels. For the deep embedding the performance results obtained with nitpick regarding counterexample finding significantly dropped in comparison to the minimal shallow embedding. The numbers reported for the deep embedding regarding the identification of formulas with no (finite) countermodels are still good though. [^11] However, some seemingly self-contradictory answers by nitpick in this setting with and without prior simplification for some problems require further consideration. Interestingly, the maximal number of 38 proofs was still obtained by sledgehammer in the deep embedding when prior simplification was applied. Future work should extend the experiments presented to more challenging test formulas $F$ and also include the further axiom schemata (and their corresponding semantic conditions) from the modal logic cube [32]. Additionally, it seems useful to replay, now using all the embeddings as provided in this paper, the metalogical cube verification experiments reported in prior work [3, 7].
@@ 9 Related work and implications for LogiKEy
The notions of deep and shallow embeddings appear to originate from the work of Boulton et al. [23], and the connections between the two have been discussed in numerous related papers, see e.g. [59, 38, 51, 60] and the references therein. However, to the best of the author’s knowledge, no previous work has succeeded in integrating deep and (various degrees of) shallow embeddings in the form presented and, in particular, in enabling mutual faithfulness proofs between them.
@@
The logic-pluralistic knowledge and reasoning methodology LogiKEy[9], see Fig. 9, leverages higher-order logic (HOL) as a unifying metalogic to enable the modeling of diverse object logics, so far, primarily via using shallow semantic embeddings. LogiKEy has been used in a range of application domains and contexts. Very different kinds of object logics have been embedded LogiKEy over the past decade, including, for example, free logic [13], modal higher-order logic [15], public announcement logic [12], conditional logic [4] and dyadic deontic logic with a preference-based semantics [49].
@@
This paper also illustrates the conceptual modifications applied in prior LogiKEy papers when transitioning from lightweight shallow embeddings of non-classical logics, such as those used to encode higher-order modal logic [11] or conditional logic [4], which consider only dependencies on possible worlds, to more heavyweight shallow embeddings, as, for example, required to embed public announcement logic [12]; in the latter work additional dependencies on evaluation domains had to be encoded. Note on the other hand, that previous work has often avoided the explicit modeling of atoms, signatures, and evaluations, which has been essential for the work presented in this paper. Moreover, the presented technique should eventually help to overcome challenges in application contexts that require the combined use of deep and shallow embeddings. The Fatio protocol for multi-agent argumentation [50] is such an example. Most importantly, the presented technique for linking deep and shallow embeddings will be very fruitful for furthering the use of LogiKEy in university-level logic courses, including the author’s courses on universal logical reasoning that have been taught for many years at FU Berlin and now at the University of Bamberg.